A speedboat travels north at 29.5 km/hr, crossing a river. The river flows east at 10.6 m/s. What is the magnitude of the boats net velocity while on the river in m/s
1
Expert's answer
2020-10-05T10:55:10-0400

The magnitude:


"v=\sqrt{v_1^2+v_2^2}\\v=\sqrt{\left(\frac{29.5}{3.6}\right)^2+10.6^2}=13.4\frac{m}{s}"